# Investigating charmed hybrid baryons via QCD sum rules

### v1

@@ Abstract
Motivated by recent interpretations of the $\eta_{1}(1855)$ and $X(2370)$ as a hybrid meson and a glueball, respectively, the possible existence of hybrid baryons has become a subject of considerable interest. In this letter, we investigate charmed hybrid baryons using the QCD sum rule method within the framework of heavy quark effective theory. We construct twenty-eight interpolating currents for charmed hybrid baryons, seven of which are employed in QCD sum rule analyses of nineteen states with quark-gluon configurations $qqcg$, $qscg$, and $sscg$ ($q=u/d$). The masses of the lowest-lying charmed hybrid baryons in the $SU(3)$ flavor $\mathbf{6}_{F}$ representation are calculated to be $M_{\Sigma_{cg}(1/2^{+})}=3.46^{+0.33}_{-0.21}~{}\rm{GeV}$, $M_{\Xi^{\prime}_{cg}(1/2^{+})}=3.60^{+0.48}_{-0.16}~{}\rm{GeV}$, and $M_{\Omega_{cg}(1/2^{+})}=3.77^{+0.31}_{-0.16}~{}\rm{GeV}$. We propose that future experiments search for these states via their $P$-wave decay channels $ND^{(*)}$, $\Lambda D^{(*)}$, and $\Xi D^{(*)}$, respectively. Such investigations would provide valuable insight into the role of gluonic excitations in hadron structure.
@@
Introduction — Quantum Chromodynamics (QCD) allows for the existence of color singlets, including conventional hadrons, multi-quark states, glueballs, and hybrids. Exploring exotic states beyond conventional hadrons is one of the most intriguing areas of research in hadronic physics, as such studies enhance our understanding of non-perturbative QCD. Gluonic degrees of freedom play a crucial role in these exotic states. The excitation of the gluonic field, analogous to the role of valence quarks, manifests in glueballs, hybrid mesons, and hybrid baryons. While substantial research has been conducted on hybrid mesons, relatively few studies have focused on hybrid baryons. Consequently, investigating hybrid baryons beyond conventional hadron configurations remains a significant and challenging topic in hadronic physics, requiring the identification of distinctive characteristics to confirm their existence. A common approach involves predicting the mass spectra and decay widths of hybrid baryons and offering specific recommendations for experimental searches.
In 2022, the BESIII collaboration observed the isoscalar state $\eta_{1}(1855)$ with the exotic quantum numbers $I^{G}J^{PC}=0^{+}1^{-+}$ BESIII:2022riz ; BESIII:2022iwi . Additionally, several collaborations IHEP-Brussels-LosAlamos-AnnecyLAPP:1988iqi ; E852:1997gvf ; E852:1998mbq ; Aoyagi:1993kn ; CrystalBarrel:1998cfz have reported three isovector states—$\pi_{1}(1400)$, $\pi_{1}(1600)$, and $\pi_{1}(2015)$—all possessing the exotic quantum numbers $I^{G}J^{PC}=1^{-}1^{-+}$. These states are regarded as strong candidates for hybrid mesons. In 2024, the BESIII collaboration also reported the $X(2370)$ as a possible lightest pseudoscalar glueball BESIII:2023wfi ; Huang:2025pyv . These observations collectively support the potential existence of hybrid baryons, although no hybrid hadron has yet been conclusively identified in experiments. Ongoing efforts to search for hybrid baryons are underway at CLAS in Hall B of Jefferson Lab Lanza:2017qel ; Burkert:2018nvj , with the first results expected in the coming years. The hybrid hadron system has also attracted considerable interest from theorists. These exotic hadrons have been studied using various theoretical frameworks, including the bag model Kusaka:1990ha ; Barnes:1977hg ; Barnes:1982fj ; Barnes:1983xx ; Duck:1983ju ; Carlson:1983xr , lattice QCD Dudek:2012ag ; Bali:1993fb ; Chen:2005mg ; Morningstar:1999rf ; Gui:2019dtm ; Michael:1985ne ; Juge:2002br ; Lacock:1996ny ; Hedditch:2005zf ; Bernard:2003jd , the large-$N_{c}$ approach Chow:1998tq , the flux-tube model Page:2004gv ; Capstick:2002wm ; Capstick:1999qq ; Page:1998qt ; Isgur:1983wj ; Qiu:2022ktc , the hybrid quark model Khosonthongkee:2014gla ; Galata:2014typ ; Li:1991sh , the constituent gluon model Horn:1977rq ; Szczepaniak:2001rg ; Guo:2007sm , and QCD sum rules Zhao:2023imq ; Kisslinger:1995yw ; Kisslinger:2003hk ; Azizi:2017xyx ; Tan:2024grd ; Lian:2024fsb , among others. Additional discussions on related developments can be found in Refs. Cimino:2024bol ; Khan:2020ahz ; Kittel:2005jm ; Page:2002mt ; Barnes:2000vn ; Chen:2022asf , which provide insights into recent progress. Notably, most studies of hybrid baryons have concentrated on systems composed entirely of either light or heavy quarks. In contrast, our focus is on singly heavy hybrid baryons containing only one heavy quark. In such systems, the heavy quark effective theory (HQET) can be effectively applied, providing a promising framework for advancing our understanding of non-perturbative QCD.
@@
In this letter, we systematically investigate charmed hybrid baryons using the QCD sum rule method within the framework of heavy quark effective theory. We construct twenty-eight interpolating currents for charmed hybrid baryons, seven of which are employed to perform sum rule analyses for nineteen states with quark-gluon configurations $qqcg$, $qscg$, and $sscg$ ($q=u/d$). We carry out comprehensive analyses of their mass spectra for quantum numbers $J^{P}=1/2^{+}$, $3/2^{+}$, and $5/2^{+}$, and determine their decay constants, which are essential for understanding their decay behaviors in future studies. In particular, we calculate the masses of the lowest-lying charmed hybrid baryons within the $SU(3)$ flavor $\mathbf{6}_{F}$ representation to be:
$M_{\Sigma_{cg}(1/2^{+})}$ $=$ $3.46^{+0.33}_{-0.21}~{}\rm{GeV}\,,$
$M_{\Xi^{\prime}_{cg}(1/2^{+})}$ $=$ $3.60^{+0.48}_{-0.16}~{}\rm{GeV}\,,$
$M_{\Omega_{cg}(1/2^{+})}$ $=$ $3.77^{+0.31}_{-0.16}~{}\rm{GeV}\,.$
@@
Categorization of charmed hybrid baryon currents — Heavy quark symmetry plays a crucial role in the study of heavy hadrons, and we apply it to systematically categorize the charmed hybrid baryon currents as follows. In the charmed hybrid baryon system, the heavy quark serves as a static color source, and the system’s properties are primarily determined by the light degrees of freedom.
Let us focus on the ground-state $Qq_{1}q_{2}g$ system, where $Q$ represents the heavy $charm$ quark, $q_{1}$ and $q_{2}$ represent the two valence light quarks ($up$, $down$, or $strange$), and $g$ represents the valence gluon. In the heavy quark limit, the total angular momentum of the system reduces to
@@
The charmed hybrid baryon currents can generally be constructed using one charm quark field $h_{v}^{a}(x)$, two light quark fields $q_{a}(x)$ and $q_{b}(x)$, and one gluon field strength tensor $G_{\mu\nu}^{n}(x)$ (or one dual gluon field strength tensor $\widetilde{G}_{\mu\nu}^{n}(x)=G^{n,\rho\sigma}(x)\times\epsilon_{\mu\nu\rho\sigma}/2$), with the color indices $a/b=1\cdots 3$ and $n=1\cdots 8$ as well as the Lorentz indices $\mu/\nu=0\cdots 3$. As illustrated in Fig. 1, we identify a total of twenty-eight charmed hybrid baryon currents, denoted as $J^{\alpha_{1}\cdots\alpha_{J-1/2}}_{F,C,j_{qq},j_{l},J^{P}}/\widetilde{J}^{\alpha_{1}\cdots\alpha_{J-1/2}}_{F,C,j_{qq},j_{l},J^{P}}$, which can be categorized into eight multiplets, denoted as $[F,C,j_{qq},j_{l},G/\widetilde{G}]$. Here, $F$ and $C$ represent the flavor and color representations of the two light quarks, while $J(\widetilde{J})$ and $G(\widetilde{G})$ indicate the presence of a gluon (or dual gluon) in the current. Their complete forms are provided in the supplemental material. We have systematically studied all these charmed hybrid baryon currents and found that the current $J_{6_{f},\bar{3}_{c},1,1,1/2^{+}}$, belonging to the doublet $[6_{f},\bar{3}_{c},1,1,G]$, couples to the lowest-lying charmed hybrid baryon. The explicit forms of this current and its spin partner for the $\Sigma_{cg}$ states are presented as follows ($q=up/down$):
@@
A point of confusion arises because these two currents exhibit a symmetric flavor structure for the two light quarks, which contrasts with the flavor structure of the lowest-lying charmed baryon $\Lambda_{c}$, but is consistent with the flavor structure of the $\Sigma_{c}$.
QCD sum rule analyses — We have systematically studied the charmed hybrid baryon currents given in Eqs. (3-4) and Eqs. (16-43) using the QCD sum rule method within the framework of heavy quark effective theory. As an example, we examine the current $J_{\Sigma_{cg},\bar{3}_{c},1,1,1/2^{+}}$ given in Eq. (3) as follows, which belongs to the doublet $[\Sigma_{cg},\bar{3}_{c},1,1,G]$. It couples to both the charmed hybrid baryon with $J^{P}=1/2^{+}$ and the one with $J^{P}=1/2^{-}$ through
@@
$=$ ${1+v\!\!\!/\over 2}\Pi_{+}(\omega)+{1-v\!\!\!/\over 2}\Pi_{-}(\omega)\,,$

### v2

@@ Abstract
We investigate charmed hybrid baryons using the QCD sum rule method within the framework of heavy quark effective theory. We construct twenty-eight interpolating currents for charmed hybrid baryons, seven of which are employed in QCD sum rule analyses of nineteen states with quark-gluon configurations qqcgqqcg, qscgqscg, and sscgsscg ($q=u/d$). The masses of the lowest-lying charmed hybrid baryons in the $SU(3)$ flavor $\mathbf{6}_{F}$ representation are calculated to be $M_{\Sigma_{cg}(1/2^{+})}=3.36^{+0.27}_{-0.26}~\rm{GeV}$, $M_{\Xi^{\prime}_{cg}(1/2^{+})}=3.59\pm 0.20~\rm{GeV}$, and $M_{\Omega_{cg}(1/2^{+})}=3.82\pm 0.21~\rm{GeV}$. We propose that future experiments search for these states via their $P$-wave decay channels $ND^{(*)}$, $\Lambda D^{(*)}$, and $\Xi D^{(*)}$, respectively. Such investigations would provide valuable insight into the role of gluonic excitations in hadron structure.
@@ I Introduction
Quantum Chromodynamics (QCD) allows for the existence of color singlets, including conventional hadrons, multi-quark states, glueballs, and hybrids. Exploring exotic states beyond conventional hadrons is one of the most intriguing areas of research in hadronic physics, as such studies enhance our understanding of non-perturbative QCD. Gluonic degrees of freedom play a crucial role in these exotic states. The excitation of the gluonic field, analogous to the role of valence quarks, manifests in glueballs, hybrid mesons, and hybrid baryons. While substantial research has been conducted on hybrid mesons, relatively few studies have focused on hybrid baryons. Consequently, investigating hybrid baryons beyond conventional hadron configurations remains a significant and challenging topic in hadronic physics, requiring the identification of distinctive characteristics to confirm their existence. A common approach involves predicting the mass spectra and decay widths of hybrid baryons and offering specific recommendations for experimental searches.
[…]
@@
In this paper we systematically investigate charmed hybrid baryons using the QCD sum rule method within the framework of heavy quark effective theory. We construct twenty-eight interpolating currents for charmed hybrid baryons, seven of which are employed to perform sum rule analyses for nineteen states with quark-gluon configurations qqcgqqcg, qscgqscg, and sscgsscg ($q=u/d$). We carry out comprehensive analyses of their mass spectra for quantum numbers $J^{P}=1/2^{+}$, $3/2^{+}$, and $5/2^{+}$, and determine their decay constants, which are essential for understanding their decay behaviors in future studies. In particular, we calculate the masses of the lowest-lying charmed hybrid baryons within the $SU(3)$ flavor $\mathbf{6}_{F}$ representation to be:
$M_{\Sigma_{cg}(1/2^{+})}$ $=$ $3.36^{+0.27}_{-0.26}~\rm{GeV}\,$,
$M_{\Xi^{\prime}_{cg}(1/2^{+})}$ $=$ $3.59\pm 0.20~\rm{GeV}\,$,
$M_{\Omega_{cg}(1/2^{+})}$ $=$ $3.82\pm 0.21~\rm{GeV}\,.$
@@
This paper is organized as follows. In Sec. II we systematically construct the interpolating currents for charmed hybrid baryons. These currents are employed in Sec. III to perform QCD sum rule analyses at leading order, and in Sec. IV to incorporate ${\mathcal{O}}(1/m_{Q})$ corrections. A summary and discussion of our findings are presented in Sec. V.
II Charmed hybrid baryon currents
In this section we systematically construct the interpolating currents for charmed hybrid baryons, where the heavy quark symmetry plays a crucial role. In the charmed hybrid baryon system, the heavy quark serves as a static color source, and the system’s properties are primarily determined by the light degrees of freedom. Let us focus on the ground-state $Qq_{1}q_{2}g$ system, where $Q$ represents the heavy charmcharm quark, $q_{1}$ and $q_{2}$ represent the two valence light quarks ($up$, downdown, or strangestrange), and $g$ represents the valence gluon. In the heavy quark limit, the total angular momentum of the system reduces to
@@ II Charmed hybrid baryon currents
The charmed hybrid baryon currents can generally be constructed using one charm quark field $h_{v}^{a}(x)$, two light quark fields $q_{a}(x)$ and $q_{b}(x)$, and one gluon field strength tensor $G_{\mu\nu}^{n}(x)$ (or one dual gluon field strength tensor $\widetilde{G}_{\mu\nu}^{n}(x)=G^{n,\rho\sigma}(x)\times\epsilon_{\mu\nu\rho\sigma}/2$), with the color indices $a/b=1\cdots 3$ and $n=1\cdots 8$ as well as the Lorentz indices $\mu/\nu=0\cdots 3$. The (dual) gluon field is a color octet, so the combination of the three quark fields should also form a color octet, with two possible combinations:
@@
As summarized in Fig. 1, we identify a total of twenty-eight charmed hybrid baryon currents, denoted as $J^{\alpha_{1}\cdots\alpha_{J-1/2}}_{F,C,j_{qq},j_{l},J^{P}}/\widetilde{J}^{\alpha_{1}\cdots\alpha_{J-1/2}}_{F,C,j_{qq},j_{l},J^{P}}$, which can be categorized into eight multiplets, denoted as $F,C,j_{qq},j_{l},G/\widetilde{G}$]. Here, $F$ and $C$ represent the flavor and color representations of the two light quarks, while J(J~)J($\widetilde${J}) and G(G~)G($\widetilde${G}) indicate the presence of a gluon (or dual gluon) in the current. The explicit expressions of these currents are given as follows:
  • The doublet $\bar{3}_{f},\bar{3}_{c},0,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $\bar{3}_{f},\bar{3}_{c},0,1,\widetilde{G}$] contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
  • The singlet $6_{f},\bar{3}_{c},1,0,G]$ contains one state with $J^{P}=1/2^{+}$, and its dual singlet $6_{f},\bar{3}_{c},1,0,\widetilde{G}$] contains another state with $J^{P}=1/2^{-}$. Their corresponding currents are
  • The doublet $6_{f},\bar{3}_{c},1,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $6_{f},\bar{3}_{c},1,1,\widetilde{G}$] contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
    $~~~~\times~~(g^{\alpha\nu}_{t}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\nu})g_{s}{\lambda_{ec}^{n}\over 2}G_{\mu\nu}^{n}h_{v}^{c}\,$,
    $~~~~\times~~(g^{\alpha\nu}_{t}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\nu})g_{s}{\lambda_{ec}^{n}\over 2}\widetilde{G}_{\mu\nu}^{n}h_{v}^{c}\,.$
  • The doublet $6_{f},\bar{3}_{c},1,2,G]$ contains two states with $J^{P}=3/2^{+}$ and $5/2^{+}$. Its dual doublet $6_{f},\bar{3}_{c},1,2,\widetilde{G}$] contains two other states with $J^{P}=3/2^{-}$ and $5/2^{-}$. Their corresponding currents are
  • The doublet $6_{f},6_{c},0,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $6_{f},6_{c},0,1,\widetilde{G}$] contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
    $~~~~~~~~~~~~~~~~~~~~\times~~\sigma_{t}^{\mu\nu}g_{s}{\lambda_{eb}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,$,
    $~~~~~~~~~~~~~~~~~~~~\times~~\sigma_{t}^{\mu\nu}g_{s}{\lambda_{eb}^{m}\over 2}\widetilde{G}_{\mu\nu}^{m}h_{v}^{c}\,$,
  • The singlet $\bar{3}_{f},6_{c},1,0,G]$ contains one state with $J^{P}=1/2^{+}$, and its dual singlet $\bar{3}_{f},6_{c},1,0,\widetilde{G}$] contains another state with $J^{P}=1/2^{-}$. Their corresponding currents are
    $~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\times~~g_{s}{\lambda_{ec}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,$,
    $~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\times~~g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\mu\nu}^{m}h_{v}^{c}\,.$
  • The doublet $\bar{3}_{f},6_{c},1,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $\bar{3}_{f},6_{c},1,1,\widetilde{G}$] contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
    $~~~~~~~~~~~\times~~\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,$,
    $~~~~~~~~~~~\times~~\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\mu\nu}^{m}h_{v}^{c}\,$,
  • The doublet $\bar{3}_{f},6_{c},1,2,G]$ contains two states with $J^{P}=3/2^{+}$ and $5/2^{+}$. Its dual doublet $\bar{3}_{f},6_{c},1,2,\widetilde{G}$] contains two other states with $J^{P}=3/2^{-}$ and $5/2^{-}$. Their corresponding currents are
    $~\times~~\Gamma_{J=2}^{\alpha\mu_{4},\mu_{1}\nu_{2}}\gamma_{t}^{\mu_{4}}\gamma_{t}^{\nu_{1}}g_{s}{\lambda_{ec}^{m}\over 2}G_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,$,
    $~\times~~\Gamma_{J=2}^{\alpha\mu_{4},\mu_{1}\nu_{2}}\gamma_{t}^{\mu_{4}}\gamma_{t}^{\nu_{1}}g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,$,
    $~~~~\times~~\Gamma_{J={5\over 2}}^{\alpha\beta,\mu_{1}\nu_{2}}\gamma_{t}^{\nu_{1}}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}G_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,$,
    $~~~~\times~~\Gamma_{J={5\over 2}}^{\alpha\beta,\mu_{1}\nu_{2}}\gamma_{t}^{\nu_{1}}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,.$
In the above expressions, $\gamma_{t}^{\mu}=\gamma^{\mu}-v\!\!\!/\penalty 50v^{\mu}$, $g_{t}^{\mu\nu}=g^{\mu\nu}-v^{\mu}v^{\nu}$, $\sigma_{t}^{\mu\nu}=\sigma^{\mu\nu}-\sigma^{\mu\alpha}v_{\alpha}v^{\nu}-\sigma^{\alpha\nu}v_{\alpha}v^{\mu}$, and $G_{\mu\nu}^{n}=\partial_{\mu}A_{\nu}^{n}-\partial_{\nu}A_{\mu}^{n}+g_{s}f^{npq}A_{p,\mu}A_{q,\nu}$; $\Gamma_{J=2}^{\alpha\beta,\mu\nu}$ and $\Gamma_{J=5/2}^{\alpha\beta,\mu\nu}$ are the $J=2$ and $J=5/2$ projection operators:
III Results at the leading order
In this section we apply the QCD sum rule method to investigate the charmed hybrid baryon currents given in Eqs. (5-18) within the framework of heavy quark effective theory. These seven currents form the four baryon multiplets $\bar{3}_{f},\bar{3}_{c},0,1,G]$, $6_{f},\bar{3}_{c},1,0,G]$, $6_{f},\bar{3}_{c},1,1,G]$, and $6_{f},\bar{3}_{c},1,2,G]$. The other currents given in Eqs. (19-32) will be investigated in our future studies. We shall find that the current $J_{6_{f},\bar{3}_{c},1,1,1/2^{+}}$, belonging to the doublet $6_{f},\bar{3}_{c},1,1,G]$, couples to the lowest-lying charmed hybrid baryon, so we use this current as a representative example. The explicit forms of this current and its spin partner for the $\Sigma_{cg}$ states are presented as follows ($q=up/down$):
@@ III Results at the leading order
The current $J_{\Sigma_{cg},\bar{3}_{c},1,1,1/2^{+}}$ couples to both the charmed hybrid baryon with $J^{P}=1/2^{+}$ and the one with $J^{P}=1/2^{-}$ through
@@
$=$ ${1+v\!\!\!/\penalty 50\over 2}\Pi_{+}(\omega)+{1-v\!\!\!/\penalty 50\over 2}\Pi_{-}(\omega)\,$,